### v1

@@ 4.1 Arrangement in the inner capsule
Figure 13 shows the arrangement in the inner capsule. All the equipments are mounted within the envelope of the inner capsule. The test area is about 830 mm in height and 500 mm in diameter, providing sufficient space for the robot to move without touching any structures. Two action cameras (GoPro HERO9) are mounted vertically and capture the robot’s motion during the sequence. Two graph papers are attached to the side walls that work as the projection screen of lasers mounted on the body 2. The trajectories of the laser lays are captured by two additional action cameras that are used to estimate the vertical releasing velocity of the robot. A cushion is placed at the bottom of the test space, which reduces damage to the robot at landing. Other electric components to control the experiment sequence are placed on the wooden platform deck above the test space.
@@ 4.2 Calibrations
The absolute angle sensors are mainly used to provide a zero-point of the joint angles. The zero-point value is measured by applying an attachment that holds the frames between bodies (Fig. 14).
@@ 5 Result
Figure 16 shows sequentially captured images by the front and the right cameras. The time stamps of the images are based on the sequence shown in Table 7. The figure shows that the robot’s attitude is reoriented after four strokes of joint actuation as described in Section 3.1. In addition, we confirmed by the video that the robot did not collide with any structures during an entire maneuver.
@@
Data of MEMS gyroscopes was successfully sampled in about Hz on average. The obtained angular velocities from all bodies are processed as shown in Fig. 17. The upper left figure is -- Euler angles of body 1, calculated by integrating the angular velocity of body 1 shown in the upper right. Initial angular velocity after the release was about deg/s. Total attitude reorientation after four joint strokes was 23.3 degrees, whereas the ideal reorientation angle of the designed maneuver is 26.1 degrees as described in Section 3.3. This gap is generated by the initial angular velocity and the difference between the commanded joint displacements and the actual (described below).
@@
The lower left figure is a history plot of joint angles. Initial angles are sampled from the absolute angle sensors, and displacements from them are calculated by integrating relative angular velocities derived by subtracting outputs of adjacent gyroscopes. Note that the initial angle of joint 1 has an offset of about degrees due to gravity torque applied before a drop. In addition, joint displacements do not precisely follow the angles shown in Table 7; e.g., the actual displacement of the first stroke is 54.8 degrees, whereas the commanded angle is 59.0 degrees. This is because the reproducibility performance of our servomotor was not high. However, the obtained result is still valid because we can evaluate the numerical model by performing numerical integration with the actual joint angle displacements.
@@
The lower right figure shows the total angular momentum of all bodies in the inertial frame, which is calculated with the obtained data and the numerical model described in Section 3.2. Ideally, the angular momentum should be conserved throughout the entire sequence, and the values before and after have agreed well. During , the conservation quality gets worse due to the influence of joint actuation. The influence of linear acceleration and vibration on MEMS gyroscopes seems to induce high-frequency noise. In contrast, low-frequency noise seems to be caused by modeling errors, such as the position of the center of mass of each body.
[…]
@@
[…]
Figure 18 shows a comparison of angular momentum between before and after the modeling error compensation; the upper figures are time-domain, and the lower are frequency-domain spectrum transformed by fast Fourier transformation, FFT [2]. The enclosed areas in the lower figure show that the compensation successfully suppressed the low-frequency noise below Hz.
@@
Figure 19 compares the frequency spectrum of angular velocity error between before and after modeling error compensation. As with angular momentum, the compensation successfully suppresses low frequency error below Hz. Figure 20 compares -- Euler angles’ error between before and after modeling error compensation. The compensation suppressed the attitude difference in the entire maneuver. Thus, we concluded that the designed modeling error compensation has successfully worked.
@@ 6.3 Error of gyroscopes
We chose cost-effective MEMS gyroscopes in this experiment because many spares must be prepared due to strong landing impacts. We can expect to use high-performance MEMS IMU for the orbit demonstration. The comparison of the major specifications between MPU-6050 and ADIS16465-2 (an example of the high-performance IMU) is shown in Table 10. Obviously, ADIS16465-2 performs much better than MPU-6050; in particular, linear acceleration sensitivity is about 10 times better. Thus, we can expect the high-frequency noise seen in Fig. 18 and 19 to be suppressed with such high-performance IMUs.
@@ 7 Conclusion
The purpose of this research was to construct an evaluation method for a numerical model for a transformable spacecraft using a drop tower facility. We introduced the design of the experiment and exhibited the results and analyses. The results showed that our numerical model replicated the robot’s motion within 1 degree of accuracy. In addition, post-experiment model corrections based on the results further improved the accuracy of the numerical solution. The performance of the hardware we used in this experiment was limited due to cost constraints, and further studies with actual flight hardware would provide a better evaluation of the actual orbit demonstration.
@@ Acknowledgements
Funding This study was supported by the Advisory Committee for Space Engineering in Japan as a strategic research group.
Competing interests This study was funded by a Grant for a strategic research group from the Advisory Committee for Space Engineering in Japan. The authors have no competing interests to declare that are relevant to the content of this article.
@@
Data availability The authors declare that the data supporting the findings of this study are available within the paper as supplementary files.
Materials availability Not applicable.
Code availability The authors declare that the source files used in the experiment are available within the paper as supplementary files.
Author contribution All authors contributed to the study conception and design. Material preparation, data collection, and analysis were performed by Tsubasa Ando, Hirona Kawahara, Shu Miyata, Naoya Uchiyama and Yuki Kubo. Design of the optimal maneuver was performed by Kazutoshi Ito. The first draft of the manuscript was written by Yuki Kubo and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.

### v2

@@ 4.1 Arrangement in the inner capsule
Figure 13 shows the arrangement in the inner capsule. All the equipment is mounted within the envelope of the inner capsule. The test area is about 830 mm in height and 500 mm in diameter, providing sufficient space for the robot to move without touching any structures. Two action cameras (GoPro HERO9) are mounted vertically and capture the robot’s motion during the sequence. Two graph papers are attached to the side walls that work as the projection screen of lasers mounted on body 2. The trajectories of the laser rays are captured by two additional action cameras that are used to estimate the vertical releasing velocity of the robot. A cushion is placed at the bottom of the test space, which reduces damage to the robot at landing. Other electric components to control the experiment sequence are placed on the wooden platform deck above the test space.
@@ 4.2 Calibrations
The absolute angle sensors are mainly used to provide a zero point for the joint angles. The zero-point value is measured by applying an attachment that holds the frames between bodies (Fig. 14).
@@ 5 Result
Figure 16 shows sequentially captured images by the front and the right cameras. The time stamps of the images are based on the sequence shown in Table 7. The figure shows that the robot’s attitude is reoriented after four strokes of joint actuation as described in Section 3.1. In addition, we confirmed by the video that the robot did not collide with any structures during the entire maneuver.
@@
Data from MEMS gyroscopes was successfully sampled in about Hz on average. The obtained angular velocities from all bodies are processed as shown in Fig. 17. The upper left figure is -- Euler angles of body 1, calculated by integrating the angular velocity of body 1 shown in the upper right. Initial angular velocity after the release was about deg/s. Total attitude reorientation after four joint strokes was 23.3 degrees, whereas the ideal reorientation angle of the designed maneuver is 26.1 degrees as described in Section 3.3. This gap is generated by the initial angular velocity and the difference between the commanded joint displacements and the actual (described below).
@@
The lower left figure is a history plot of joint angles. Initial angles are sampled from the absolute angle sensors, and displacements from them are calculated by integrating relative angular velocities derived by subtracting outputs of adjacent gyroscopes. Note that the initial angle of joint 1 has an offset of about degrees due to the gravity torque applied before a drop. In addition, joint displacements do not precisely follow the angles shown in Table 7; e.g., the actual displacement of the first stroke is 54.8 degrees, whereas the commanded angle is 59.0 degrees. This is because the reproducibility performance of our servomotor was not high. However, the obtained result is still valid because we can evaluate the numerical model by performing numerical integration with the actual joint angle displacements.
@@
The lower right figure shows the total angular momentum of all bodies in the inertial frame, which is calculated with the obtained data and the numerical model described in Section 3.2. Ideally, the angular momentum should be conserved throughout the entire sequence, and the values before and after agree well. During , the conservation quality gets worse due to the influence of joint actuation. The influence of linear acceleration and vibration on MEMS gyroscopes seems to induce high-frequency noise. In contrast, low-frequency noise seems to be caused by modeling errors, such as the position of the center of mass of each body.
The latter error can be suppressed by compensating for the modeling error of the center of mass of each body. Thus, we solved an optimization problem defined in Eq. (5) to obtain a proper offset of the center of mass. Optimized variables are , position offsets of center of mass in the local frame of -th body . denotes total angular momentum of the robot at and the bar signifies moving average of local -points. Thus, the objective function is the sum of the norm of the difference between the moving-averaged angular momentum at and for all sampling points. The moving average works as a simple low-pass filter, and the width of the averaging window is determined so that it satisfies the appropriate cutoff frequency. In this case, is set; it corresponds to cutoff frequency Hz in sampling frequency Hz [33]. are magnitudes of maximum offsets. We set mm for and mm because the body 4 does not move with respect to the body 3 in this maneuver. Table 9 shows the optimization result.
@@
Figure 18 shows a comparison of angular momentum between before and after the modeling error compensation; the upper figures are time-domain, and the lower are frequency-domain spectra transformed by fast Fourier transformation, FFT [34]. The enclosed areas in the lower figure show that the compensation successfully suppressed the low-frequency noise below Hz.
@@
Figure 19 compares the frequency spectrum of angular velocity error between before and after modeling error compensation. As with angular momentum, the compensation successfully suppresses low-frequency error below Hz. Figure 20 compares -- Euler angles’ error between before and after modeling error compensation. The compensation suppressed the attitude difference in the entire maneuver. Before the compensation, the final attitude error between the numerical simulation and the experiment data in -- Euler angle is deg, and its norm is 0.31 deg. On the other hand, attitude error after the compensation is deg, and its norm is 0.24 deg, which corresponds to 30 % error reduction. The performance in this experiment is evaluated in Section 6.1.
@@ 6 Discussion
6.1 Expected on-orbit performance
From the results obtained in the drop tower test, we can evaluate the on-orbit performance of a spacecraft. As a potential example of the spacecraft, we consider the 3U-sized transformable CubeSat designed by the authors shown in Fig. 21.
It is designed to have similar dimensions and joint configurations to the robot used in the drop tower test. However, the drop-tower robot is designed for high joint actuation speed and acceleration to complete the maneuver in the limited micro-gravity duration. In contrast, the on-orbit spacecraft actuates its joint much more slowly to maximize its torque with a smaller actuator. The joint speed performance of the drop-tower robot can be estimated by the lower left graph in Fig. 17 as deg/s. On the other hand, the designed average joint speed performance of the 3U CubeSat is deg/s.
In terms of reorientation maneuvers, the on-orbit CubeSat can perform various maneuvers because there is no time limit to complete the maneuver, unlike the drop tower. Figure 22 shows a distribution of nonholonomic attitude reorientation of the 3U transformable CubeSat where joint angles are discretized in a 5-degree interval. Joint actuation sequence is a four-stroke alternating maneuver similar to Fig. 8 and all combinations of joint pairs are sampled to construct the distribution. The color bar shows the reorientation efficiency, which is calculated as the magnitude of the reorientation angle divided by the total joint actuation angle. The efficiency of the maneuver designed in the drop tower test was (cf. Section 3.3). The optimal maneuver in the distribution is the point shown in the yellow star in Fig. 22, which corresponds to 39.31 degrees reorientation with actuation of deg and deg in this order (vice versa in the third and fourth stroke). The efficiency of the maneuver is , which is slightly better than the maneuver in the drop tower test. Thus, the expected average reorientation speed of the 3U transformable CubeSat can be estimated as deg/s.
Now, we can evaluate the reorientation performance of the 3U transformable CubeSat. The conclusion in Section 5 was that the numerical model before the modeling error compensation can design the attitude maneuver with 0.314 degree accuracy for the 23.3 degree reorientation, whereas the model after the compensation can design with 0.237 degree accuracy. When designing a 90-degree reorientation on-orbit, the accuracy is expected to be deg with the model without the compensation, while deg with the model with the compensation. Precise pointing cannot be achieved solely by the nonholonomic attitude control, but coarse and rapid attitude reorientation can be achieved. This accuracy seems to be the worst-case value because we can expect more precise modeling error compensation with rich on-orbit flight data as described later in Section 6.3. The maneuver time, assuming the average reorientation speed, is expected to be sec. According to previous research, the average slew rate in recent Earth observation satellite missions is around 0.1-1 deg/s, and a 3-10 deg/s slew rate is classified as a highly challenging level of agility [35][36]. In conclusion, the drop tower test confirmed that the 3U transformable CubeSat system can overcome conventional maneuverability with acceptable accuracy. The reliability of this performance estimation will be examined in the future on-orbit flight experiment with the 3U transformable CubeSat.
@@ 6.4 Error of gyroscopes
We chose cost-effective MEMS gyroscopes in this experiment because many spares must be prepared due to strong landing impacts. We can expect to use a high-performance MEMS IMU for the orbit demonstration. The comparison of the major specifications between MPU-6050 and ADIS16465-2 (an example of the high-performance IMU) is shown in Table 10. ADIS16465-2 performs much better than MPU-6050; in particular, linear acceleration sensitivity is about 10 times better. Thus, we can expect the high-frequency noise seen in Fig. 18 and 19 to be suppressed with such high-performance IMUs.
@@ 7 Conclusion
The purpose of this research was to develop an evaluation method for a numerical model of a transformable spacecraft, utilizing a drop tower facility. We introduced the design of the experiment and exhibited the results and analyses. The results showed that our numerical model replicated the robot’s motion within 1 degree of accuracy. In addition, post-experiment model corrections based on the results further improved the accuracy of the numerical solution. The performance of the hardware we used in this experiment was limited due to cost constraints, and further studies with actual flight hardware would provide a better evaluation of the actual orbit demonstration.
@@ Acknowledgements
Funding This study was supported by the Advisory Committee for Space Engineering in Japan as a strategic research group. The funding supporter has not been involved in the preparation of this paper.
Competing interests The authors have no competing interests to declare that are relevant to the content of this article.
@@
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used DeepL and Grammarly in order to improve readability of the text. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.